Estimate the pH of natural rainwater in a heavily polluted industrial area, assuming that the only...

  1. Estimate the pH of natural rainwater in a heavily polluted industrial area, assuming that the only substance affecting the it is the adsorption of CO2 from the atmosphere. Assume that the concentration of CO2 is 360 ppm and the temperature and pressure are 25°C and 1 atm, respectively. Other parameters: KH=29.97 atm-L/mol, wind speed = 25 m/s, UV energy density =2.25 J/cm2, and benzene concentration = 14 ppb. (material will be discussed in class on 1/31/2020)
Answer #1

Hi
Concentration of CO2 in the air is given as 360 ppm.
Since nothing is mentioned in the case of pollutants we take ppmv and thus we can write.
mole fraction of the CO2 =360*10^-6

Using Dalton's law of partial pressure we can write that.
Pa=Pt*Xa
Pa is the partial pressure of component a.
Pt is total pressure and xa is the mole fraction of the component a.
Total pressure can be assumed to be 1 atm.
Pa=1*360*10^-6 atm

Using Henry law.
Pa=KH*[A]
360*10^-6=29.97*[A]
[A]=1.2*10^-5 mole/L
[A] means concentration of A

We have A as CO2 here.
We can assume that CO2 will be in the equilibrium as per the following relation .

H2CO3- 2H+ +CO3-2

We can assume that all H2CO3 is getting dissociated since nothing is mentioned.


[H+ ]= 2*1.2*10^-5 =2.4*10^-5 M
one mole of H2CO3 will give 2 mole of H+.

pH=-log[H+]
=-log(2.4*10^-5)
=4.61

Hope this helps

Thanks